# Distributed Hybrid Sketching for $\ell_{2}$-Embeddings

### v1

@@ III Sketching Overview
We denote $\mathbb{N}_{m}\coloneqq\{1,2,\ldots,m\}$, and $X_{\{m\}}=\{X_{i}\}_{i=1}^{m}$; where $X$ can be replaced by any variable. We consider random square projection matrices; which are represented by $\mathbf{P}$. Sampling matrices are denoted by $\mathbf{\Omega}$, where each row has a single nonzero entry. The uniform sampling distribution $\{1/N,\ldots,1/N\}$ is denoted by $\mathcal{U}_{N}$, and by $\tilde{\mathcal{U}}_{N}$ we represent $\mathcal{U}_{N}$’s approximation through our approach. The index set of $\mathbf{A}$’s rows is denoted by $\mathcal{I}$, i.e. $\mathcal{I}=\mathbb{N}_{N}$. The index set of $\mathbf{A}_{i}$ is denoted by $\mathcal{I}_{i}$ for each $i\in\mathbb{N}_{k}$, i.e. $\mathcal{I}_{i}=\{(i-1)n+1,\ldots,in\}$ and $\mathcal{I}=\bigsqcup_{\iota=1}^{k}\mathcal{I}_{\iota}$. Similarly, by $\mathcal{S}$ we denote the index multiset of $\mathbf{A}$’s sampled rows, and $\mathcal{S}_{i}$ the index multiset of $\mathbf{A}_{i}$’s sampled rows for each $i$. By $\mathbf{e}_{j}$ we denote the $j^{th}$ standard basis vector. The restriction of a matrix $\mathbf{M}$ to the entries of $\mathcal{I}_{\iota}$ is represented by $\mathbf{M}\big{|}_{\mathcal{I}_{\iota}}$, e.g. $\mathbf{I}_{n}\big{|}_{(\{1,2\})}=(\mathbf{e}_{1}\mathbf{e}_{1}^{\top}+\mathbf{e}_{2}\mathbf{e}_{2}^{\top})\in\mathbb{R}^{n\times n}$ The $j^{th}$ row of matrix $\mathbf{M}$ is represented by $\mathbf{M}_{(j)}$. We abbreviate $(1-\epsilon)\cdot b\leqslant a\leqslant(1+\epsilon)\cdot b$ to $a\leqslant_{\epsilon}b$, and the complementary event of $|a-b|>\epsilon\cdot b$ to $a>_{\epsilon}b$. We use $\wedge$” and $\sim$” (over-scripts) to denote corresponding quantities of $\mathbf{A}$ in the global $\widehat{\mathbf{A}}$ and hybrid $\widetilde{\mathbf{A}}$ sketches respectively.
@@ III-A Sketching through Leverage Scores
Many sampling algorithms select data points according to the data’s leverage scores [54, 55]. The leverage scores of $\mathbf{A}$ measure the extent to which the vectors of its orthonormal basis $\mathbf{U}$ are correlated with the standard basis, and define the key structural non-uniformity that must be dealt with when developing fast randomized matrix algorithms; as they characterize the importance of the data points. Leverage scores are defined as $\ell_{j}\coloneqq\|\mathbf{U}_{(j)}\|_{2}^{2}$; and are agnostic to any particular basis, as they are equal to the diagonal entries of the projection matrix $P_{\mathbf{A}}=\mathbf{A}\mathbf{A}^{\dagger}=\mathbf{U}\mathbf{U}^{\top}$. The normalized leverage scores of $\mathbf{A}$ for each $j\in\mathbb{N}_{N}$ are
@@
and $\pi_{\{N\}}$ form a sampling probability distribution; as $\sum_{\iota=1}^{N}\pi_{\iota}=1$ and $\pi_{j}\geqslant 0$ for all $j$. This induced distribution has been proven useful in linear regression [55, 3, 43, 35], as well as a plethora of other applications [53, 8, 4, 56].
@@
and define the “local coherence” of each data block as $\gamma_{i}\coloneqq\max_{j\in\mathcal{I}_{i}}\big{\{}\ell_{j}\big{\}}$. We denote the sum of each block’s corresponding leverage scores by $\mathscr{L}_{i}\coloneqq\sum_{j\in\mathcal{I}_{i}}\ell_{j}$. It is worth noting that in our setting, the closer $\mathscr{L}_{i}$ and $\gamma_{i}$ are to $d/k$ and $d/N$ respectively, the more homogeneous the local data blocks are; when they considered as a global dataset. If $\gamma_{i}=d/N$, $\mathbf{U}_{i}$ is aligned with the standard basis.
Next, we recall the leverage score sampling $\ell_{2}$-s.e. sketch; which is used to obtain ${\grave{\mathbf{A}}}\coloneqq\grave{\mathbf{S}}\mathbf{A}$. Given $\mathbf{A}$, we sample $R>d$ rows with replacement (w.r.) from $\mathbf{A}$ according to $\pi_{\{N\}}$. If at trial $j$ the row $i_{j}$ was sampled; we rescale it $1/\sqrt{R\pi_{i_{j}}}$, and set ${\grave{\mathbf{A}}}_{(j)}=\mathbf{A}_{(i_{j})}\big{/}\sqrt{R\pi_{i_{j}}}$. It is clear that here $\grave{\mathbf{S}}\in\mathbb{R}^{R\times N}$ is simply a sampling and rescaling matrix, i.e. $\grave{\mathbf{S}}_{j}=\mathbf{e}_{i_{j}}^{\top}\big{/}\sqrt{R\pi_{i_{j}}}$.
In many cases, estimating the leverage scores is preferred, as computing them exactly requires $\mathcal{O}(Nd^{2})$ time which is excessive. We can instead use accurate approximate scores ~$\tilde{\ell}_{\{N\}}$ which can be computed in $\mathcal{O}(Nd\log N)$ time [55]. The estimates are “close” in the following sense: ~$\tilde{\ell}_{i}\geqslant\beta\ell_{i}$ for all $i$, where $\beta\in(0,1]$ is a misestimation factor. The only difference in sampling according to ~$\tilde{\ell}_{\{N\}}$, is that we need to oversample by a factor of $1/\beta$ to get the same theoretical guarantee. The $\ell_{2}$-s.e. result of $\grave{\mathbf{S}}$ is presented next [3, 35, 22].
@@
[…]
@@ III-B Oblivious Subspace Embeddings
A drawback of directly applying leverage score sampling locally in hope of obtaining a global sketch by aggregating the local sketches, is that the local datasets may be highly heterogeneous, and the sampling performed locally may not be representative of the global leverage score sampling distribution. To alleviate this issue, we resort to OSEs, which exploit random projections and/or uniform sampling. Two prime examples of OSEs, are the Gaussian sketch and the SRHT. It is also worth noting that utilizing random Gaussian and Rademacher random matrices in data compression has close ties to the Johnson-Lindenstrauss lemma [57, 58, 59], which predates the study of RandNLA. We note that Rademacher sketching matrices are also referred to as “Sub-Gaussian sketches” in the literature, since the tail of Rademacher random variables follow a sub-Gaussian decay, i.e. its tail probabilities decay at least as fast as those of a normal distribution.
@@
The idea behind our approach is that the local projections will “flatten” the leverage scores in $\mathbf{A}$ of their local blocks; i.e. $\pi_{j}\approx\mathscr{L}_{i}/(nd)$ for each $j\in\mathcal{I}_{i}$ and every $i\in\mathbb{N}_{k}$. By then locally performing uniform row sampling on $\mathbf{P}_{i}\mathbf{A}_{i}$; we get a close to uniform sampling across all the projected blocks.
@@
Since the resulting sketching matrix $\hat{\mathbf{S}}=\mathbf{\Pi}_{2}\mathbf{\Pi}_{1}$ is composed of two distinct linear transformations, we cannot hope to get an improvement individually on one of the two characteristics discussed above over all sketching techniques, though we do get a better balance on the two; when compared to the individual sketching approaches. By the aforementioned selection of $\mathbf{\Pi}_{1}$ and $\mathbf{\Pi}_{2}$; presented in the last row of Table I, we therefore obtain an interpolation between the two extreme cases of OSEs. As we elaborate on in Section V, a primary objective for applying a dense $\mathbf{\Pi}_{2}$ in our distributed setting, is to obtain linear combinations of all data points in the final sketch, by a recombination of the local sketches. In this scenario, $\mathbf{\Pi}_{1}$ can be interpreted as a sparse sketching matrix (8).
Comparison to Importance Sampling: From Table I, we note that in worst cases the $\log(d/\delta)$ factor in the leverage score sampling sketch which is non-oblivious, implies that leverage score sketches have a theoretically higher sample complexity $R$. However, this is a worst case scenario, which occurs when $\gamma(\mathbf{A})\approx d/N$; i.e. $\mathbf{A}$ has low coherence and the leverage scores across the rows are close to uniform. This is also the assumption that other works make to simplify their analyses, which we bypass through local projections; as was mentioned in Sec. I. Moreover, the effect of leverage score sampling is beneficial when $\gamma(\mathbf{A})$ is high, meaning that a small subset of rows is of greater importance than the rest, and require a lower sample complexity in practice to maintain the structure of $\mathbf{A}$. High leverage rows are critical for preserving the span of the matrix’s principal components, allowing leverage score sketches to capture essential information with fewer samples than a Gaussian sketch, especially when $\mathbf{A}$ is incoherent. This point is captured in the case of uniform sampling for an $\ell_{2}$-s.e. in the $5^{th}$ row of Table I, where the embedding dimension is proportional to the coherence of the subspace. In such cases, leverage score sketches can provide accurate results with fewer samples than the required $R$ indicated above, due to the fact that they then perform non-uniform targeted sampling, which often reduces the effective sample size required in practice. The Gaussian and Rademacher sketches which are OSEs, drop this dependency on $\gamma(\mathbf{A})$, thus uniform sampling suffices; and their analyses do not need to consider a worst case scenario.
@@ IV Distributed Local Sketching
In this section, we discuss the details of our distributed sketching scheme (Algorithm 1). In the setup of Figure 1, the $i^{th}$ node applies a random projection matrix $\mathbf{P}_{i}\in\mathbb{R}^{n\times n}$ which is generated locally; to flatten the corresponding leverage scores. As we will see, the flattening here is with respect to (w.r.t.) $\mathbf{U}_{i}$; i.e. w.h.p. $\ell_{j}\approx\mathscr{L}_{i}/(nd)$ for each $j\in\mathcal{I}_{i}$. This is the cost we pay for performing local sketching. Nonetheless, we show in Figures 2 that the flattening degrades gracefully as $k$ increases; even for global datasets with highly non-uniform leverage scores. To partially circumvent this concern if our approach is to be performed by a single user or centrally administered by the coordinator, a random permutation can be applied on the rows of $\mathbf{A}$ before the partitioning takes place.
After locally applying $\mathbf{P}_{i}$, the nodes randomly sample $r=R/k$ rows from $\mathbf{P}_{i}\mathbf{A}_{i}$ which they rescale by $\sqrt{n/r}=\sqrt{N/R}$ and aggregate through $\mathbf{\Omega}_{i}$; in order to obtain the local sketches
@@
Then, each node communicates the resulting sketch to the coordinator; who aggregates them.
Input: Effective local dimension $r$triangleright $R>d$ and $R=rk$
@@
       1)\ 1)1 ) Generate a random $\mathbf{P}_{i}\in\mathbb{R}^{n\times n}$triangleright $\mathbb{E}\left[\mathbf{P}_{i}^{\top}\mathbf{P}_{i}\right]=\mathbf{I}_{n}$
       2)\ 2)2 ) Uniformly sample $r$ rows from $\mathbf{P}_{i}\mathbf{A}_{i}$, through $\mathbf{\Omega}_{i}$
       3)\ 3)3 ) Deliver $\sqrt{\frac{n}{r}}\left(\mathbf{\Omega}_{i}\mathbf{P}_{i}\right)\mathbf{A}_{i}\eqqcolon\mathbf{S}_{i}\mathbf{A}_{i}$ to the coordinator
       note: 1)1)1 ) and 2)2)2 ) can be performed simultaneously by generating $\mathbf{S}_{i}\in\mathbb{R}^{r\times n}$, to reduce the local computations
@@ IV-A Analysis of our approach
${\overbrace{\begin{pmatrix}\mathbf{S}_{1}&&\\ &\ddots&\\ &&\mathbf{S}_{k}\end{pmatrix}}^{\widehat{\mathbf{S}}\in\mathbb{R}^{R\times N}}=\sqrt{\frac{N}{R}}\cdot\overbrace{\begin{pmatrix}\mathbf{\Omega}_{1}&&\\ &\ddots&\\ &&\mathbf{\Omega}_{k}\end{pmatrix}}^{\hat{\mathbf{\Omega}}\in\{0,1\}^{R\times N}}\cdot\overbrace{\begin{pmatrix}\mathbf{P}_{1}&&\\ &\ddots&\\ &&\mathbf{P}_{k}\end{pmatrix}}^{\hat{\mathbf{P}}\in\mathbb{R}^{N\times N}}}$ (8)
@@
where $\hat{\mathbf{U}}_{i}=\big{(}\sqrt{n/r}\cdot\mathbf{P}_{i}\big{)}\mathbf{U}_{i}$ for each $i\in\mathbb{N}_{k}$. We denote the leverage scores of the projected matrix $\hat{\mathbf{P}}\mathbf{A}$ by ^$\hat{\ell}_{j}=\|\hat{\mathbf{U}}_{(j)}\|_{2}^{2}$.
Next, we show that the leverage scores of each $\hat{\mathbf{U}}_{i}$ are flattened w.r.t. $\mathscr{L}_{i}$ by the local projection. For our analysis, we assume that $\mathbf{P}_{i}$ is a random unitary matrix, drawn from an arbitrary large finite subset $\tilde{O}_{n}(\mathbb{R})$ of the set orthonormal matrices $O_{n}(\mathbb{R})$ of size $n\times n$. Such sketching techniques; in which we first project with a unitary matrix and and then uniformly sample rows from the transformed matrix, are referred to as “Randomized Orthonormal Systems” [49]. The SRHT is a special case and the most popular, due to to the fact the the Hadamard matrix is well structured and can be applied efficiently. In practice, normalized Gaussian and Rademacher matrices also widely used; as they satisfy $\mathbb{E}\left[\mathbf{G}^{\top}\mathbf{G}\right]=\mathbf{I}_{N}$. Analogous results can also be derived for these options of $\mathbf{P}_{i}$.
@@
For a fixed $i\in\mathbb{N}_{k}$ and $\xi>0$, the normalized (w.r.t. $\hat{\mathbf{U}}$) leverage scores {¯$\{\bar{\ell}_{j}\}_{j\in\mathcal{I}_{i}}$ corresponding to $\hat{\mathbf{U}}_{i}$ satisfy
@@
In Figures 2, we show numerically the flattening of the normalized leverage scores of a random $\mathbf{A}\in\mathbb{R}^{2000\times 40}$ following a $t$-distribution, which scores were highly non-uniform. In this experiments, $\mathbf{P}_{\{k\}}$ were random Gaussian matrices. As noted previously, random Gaussian matrices are good surrogates for unitary random matrices, and are widely used in practice as they are approximately orthogonal. Furthermore, we observe that the flattening degrades gracefully as $k$ increases, and if a random permutation on $\mathbf{A}$’s rows is applied before $\hat{\mathbf{P}}$, we have slightly better results for each $k$. Analogous simulation results were observed when the $\mathbf{P}_{i}$’s were randomized Hadamard transforms, random unitary, and Rademacher random matrices.
@@
Before we present our main result, we need to show that $\mathbf{\Omega}$ is close to being a uniform sampling (w.r.) matrix of $R$ out of $N$ rows. In Proposition 2, we show this by applying Chebyshev’s inequality to the balls into bins problem. Specifically, it shows that w.h.p. the sampling of $\mathcal{S}$ in $\mathcal{I}$ is close to a sampling of $\bigcup_{i=1}^{k}\mathcal{S}_{i}$, i.e. the cardinality of sampled indices of $\mathcal{S}$ that lie in any $\mathcal{I}_{i}$ is not far from #$\#\mathcal{S}_{i}=r$. The important factors to note are that all sampling trials in both scenarios are uniform, identical, independent and with replacement.
@@
Partition the sampled index set $\mathcal{S}$ into ordered partitions $S_{i}$ of $\mathcal{I}$; according to $\mathcal{I}_{\{k\}}$, i.e. $S_{i}=\mathcal{S}\bigcap\big{(}\bigcup_{l=1}^{R}\mathcal{I}_{i}\big{)}$. Then, for any $i\in\mathbb{N}_{k}$: $\Pr\big{[}|\#S_{i}-r|\geqslant 10\big{]}\leqslant 1/100$.
@@
Let $\mathbf{P}_{\{k\}}$ of Algorithm 1 be random unitary matrices, and $\hat{\beta}=\frac{k}{d}\cdot\min_{i\in\mathbb{N}_{k}}\left\{\mathscr{L}_{i}\right\}$. Then, for $\delta>0$ and $R=\Theta\left(d\log{(2d/\delta)}/(\hat{\beta}\epsilon^{2})\right)$, the sketching matrix $\widehat{\mathbf{S}}$ of the global $\mathbf{A}$, satisfies (5) with high probability.
@@
Consider $\frac{1}{\sqrt{r}}\mathbf{G}_{\{k\}}$ rescaled random Gaussian matrices, and perform Gram-Schmidt to each projection to obtain $\hat{\mathbf{P}}_{\{k\}}$. Then, for $\delta>0$ and $R=\Theta\left(d\log{(2d/\delta)}/(\hat{\beta}\epsilon^{2})\right)$, the sketching matrix $\widehat{\mathbf{S}}$ of the global $\mathbf{A}$, w.h.p. satisfies (5).

### v2

@@ III Sketching Overview
We denote $\mathbb{N}_{m}\coloneqq\{1,2,\ldots,m\}$, and $X_{\{m\}}=\{X_{i}\}_{i=1}^{m}$ where $X$ can be replaced by any variable. We consider random square projection matrices, which are represented by $\mathbf{P}$. Sampling matrices are denoted by $\mathbf{\Omega}$, where each row has a single nonzero entry. The uniform sampling distribution $\{1/N,\ldots,1/N\}$ is denoted by $\mathcal{U}_{N}$, and by $\tilde{\mathcal{U}}_{N}$ we represent $\mathcal{U}_{N}$’s approximation through our approach. The index set of $\mathbf{A}$’s rows is denoted by $\mathcal{I}$, i.e. $\mathcal{I}=\mathbb{N}_{N}$. The index set of $\mathbf{A}_{i}$ is denoted by $\mathcal{I}_{i}$ for each $i\in\mathbb{N}_{k}$, i.e. $\mathcal{I}_{i}=\{(i-1)n+1,\ldots,in\}$ and $\mathcal{I}=\bigsqcup_{\iota=1}^{k}\mathcal{I}_{\iota}$. Similarly, by $\mathcal{S}$ we denote the index multiset of $\mathbf{A}$’s sampled rows, and $\mathcal{S}_{i}$ the index multiset of $\mathbf{A}_{i}$’s sampled rows for each $i$. By $\mathbf{e}_{j}$ we denote the $j^{th}$ standard basis vector. The restriction of a matrix $\mathbf{M}$ to the entries of $\mathcal{I}_{\iota}$ is represented by $\mathbf{M}\big|_{\mathcal{I}_{\iota}}$, e.g. $\mathbf{I}_{n}\big|_{(\{1,2\})}=(\mathbf{e}_{1}\mathbf{e}_{1}^{\top}+\mathbf{e}_{2}\mathbf{e}_{2}^{\top})\in\mathbb{R}^{n\times n}$ . The $j^{th}$ row of matrix $\mathbf{M}$ is represented by $\mathbf{M}_{(j)}$. We abbreviate $(1-\epsilon)\cdot b\leqslant a\leqslant(1+\epsilon)\cdot b$ to $a\leqslant_{\epsilon}b$, and the complementary event of $|a-b|>\epsilon\cdot b$ to $a>_{\epsilon}b$. We use $\wedge$” and $\sim$” (over-scripts) to denote corresponding quantities of $\mathbf{A}$ in the global $\widehat{\mathbf{A}}$ and hybrid $\widetilde{\mathbf{A}}$ sketches respectively.
@@ III-A Sketching through Leverage Scores
Many sampling algorithms select data points according to the data’s leverage scores [53, 54]. The leverage scores of $\mathbf{A}$ measure the extent to which the vectors of its orthonormal basis $\mathbf{U}$ are correlated with the standard basis, and define the key structural non-uniformity that must be dealt with when developing fast randomized matrix algorithms, as they characterize the importance of the data points. Leverage scores are defined as $\ell_{j}\coloneqq\|\mathbf{U}_{(j)}\|_{2}^{2}$ and are agnostic to any particular basis, as they are equal to the diagonal entries of the projection matrix $P_{\mathbf{A}}=\mathbf{A}\mathbf{A}^{\dagger}=\mathbf{U}\mathbf{U}^{\top}$. The normalized leverage scores of $\mathbf{A}$ for each $j\in\mathbb{N}_{N}$ are
@@
and $\pi_{\{N\}}$ form a sampling probability distribution, as $\sum_{\iota=1}^{N}\pi_{\iota}=1$ and $\pi_{j}\geqslant 0$ for all $j$. This induced distribution has been proven useful in linear regression [54, 3, 43, 35], as well as a plethora of other applications [55, 8, 4, 56].
@@
and define the “local coherence” of each data block as $\gamma_{i}\coloneqq\max_{j\in\mathcal{I}_{i}}\big\{\ell_{j}\big\}$. We denote the sum of each block’s corresponding leverage scores by $\mathscr{L}_{i}\coloneqq\sum_{j\in\mathcal{I}_{i}}\ell_{j}$. It is worth noting that in our setting, the closer $\mathscr{L}_{i}$ and $\gamma_{i}$ are to $d/k$ and $d/N$ respectively, the more homogeneous the local data blocks are when they considered as a global dataset. If $\gamma_{i}=d/N$, $\mathbf{U}_{i}$ is aligned with the standard basis.
Next, we recall the leverage score sampling $\ell_{2}$-s.e. sketch, which is used to obtain ${\grave{\mathbf{A}}}\coloneqq\grave{\mathbf{S}}\mathbf{A}$. Given $\mathbf{A}$, we sample $R>d$ rows with replacement (w.r.) from $\mathbf{A}$ according to $\pi_{\{N\}}$. If at trial $j$ the row $i_{j}$ was sampled we rescale it by $1/\sqrt{R\pi_{i_{j}}}$, and set ${\grave{\mathbf{A}}}_{(j)}=\mathbf{A}_{(i_{j})}\big/\sqrt{R\pi_{i_{j}}}$. It is clear that here $\grave{\mathbf{S}}\in\mathbb{R}^{R\times N}$ is simply a sampling and rescaling matrix, i.e. $\grave{\mathbf{S}}_{j}=\mathbf{e}_{i_{j}}^{\top}\big/\sqrt{R\pi_{i_{j}}}$.
In many cases, estimating the leverage scores is preferred, as computing them exactly requires $\mathcal{O}(Nd^{2})$ time which is excessive. We can instead use accurate approximate scores $\tilde{\ell}_{\{N\}}$ which can be computed in $\mathcal{O}(Nd\log N)$ time [54]. The estimates are “close” in the following sense: $\tilde{\ell}_{i}\geqslant\beta\ell_{i}$ for all $i$, where $\beta\in(0,1]$ is a misestimation factor. The only difference in sampling according to $\tilde{\ell}_{\{N\}}$, is that we need to oversample by a factor of $1/\beta$ to get the same theoretical guarantee. The $\ell_{2}$-s.e. result of $\grave{\mathbf{S}}$ is presented next [3, 35, 22].
@@
The leverage score sketching matrix $\grave{\mathbf{S}}$ is a $\ell_{2}$-s.e of $\mathbf{A}$. Specifically, for $\delta>0$ and $R=\Theta\left(d\log{(2d/\delta)}/(\beta\epsilon^{2})\right)$, the identity of (5) is satisfied with probability at least $1-\delta$.
@@ III-B Oblivious Subspace Embeddings
A drawback of directly applying leverage score sampling locally in hope of obtaining a global sketch by aggregating the local sketches, is that the local datasets may be highly heterogeneous, and the sampling performed locally may not be representative of the global leverage score sampling distribution. To alleviate this issue, we resort to OSEs, which exploit random projections and/or uniform sampling.
Definition 2.
A Random matrix $\mathbf{\Pi}\in\mathbb{R}^{R\times N}$ is an $(\epsilon,\delta,d)$-oblivious subspace embedding (OSE) if for any $d$-dimensional subspace $T=\mathsf{im}(\mathbf{U})\subseteq\mathbb{R}^{N}$, it satisfies (5) for $\mathbf{S}\leftarrow\mathbf{\Pi}$.
Two prime examples of OSEs, are the Gaussian sketch and the SRHT. It is also worth noting that utilizing random Gaussian and Rademacher random matrices in data compression has close ties to the Johnson-Lindenstrauss lemma [57, 58, 59], which predates the study of RandNLA. We note that Rademacher sketching matrices are also referred to as “Sub-Gaussian sketches” in the literature, since the tail of Rademacher random variables follow a sub-Gaussian decay, i.e. its tail probabilities decay at least as fast as those of a normal distribution.
@@
The idea behind our approach is that the local projections will “flatten” the leverage scores in $\mathbf{A}$ of their local blocks, i.e. $\pi_{j}\approx\mathscr{L}_{i}/(nd)$ for each $j\in\mathcal{I}_{i}$ and every $i\in\mathbb{N}_{k}$. By then locally performing uniform row sampling on $\mathbf{P}_{i}\mathbf{A}_{i}$, we get a close to uniform sampling across all the projected blocks.
@@
Since the resulting sketching matrix $\hat{\mathbf{S}}=\mathbf{\Pi}_{2}\mathbf{\Pi}_{1}$ is composed of two distinct linear transformations, we cannot hope to get an improvement individually on one of the two characteristics discussed above over all sketching techniques. We do however get a better balance on the two, when compared to the individual sketching approaches. By the aforementioned selection of $\mathbf{\Pi}_{1}$ and $\mathbf{\Pi}_{2}$ presented in the last row of Table I, we therefore obtain an interpolation between the two extreme cases of OSEs. As we elaborate on in Sec. V, a primary objective for applying a dense $\mathbf{\Pi}_{2}$ in our distributed setting, is to obtain linear combinations of all data points in the final sketch, by a recombination of the local sketches. In this scenario, $\mathbf{\Pi}_{1}$ can be interpreted as a sparse sketching matrix (8).
Comparison to Importance Sampling: From Table I, we note that in worst cases the $\log(d/\delta)$ factor in the leverage score sampling sketch which is non-oblivious, implies that leverage score sketches have a theoretically higher sample complexity $R$. However, this is a worst case scenario, which occurs when $\gamma(\mathbf{A})\approx d/N$; i.e. $\mathbf{A}$ has low coherence and the leverage scores across the rows are close to uniform. This is also the assumption that other works make to simplify their analyses, which we bypass through local projections, as was mentioned in Sec. I. Moreover, the effect of leverage score sampling is beneficial when $\gamma(\mathbf{A})$ is high, meaning that a small subset of rows is of greater importance than the rest, and require a lower sample complexity in practice to maintain the structure of $\mathbf{A}$. High leverage rows are critical for preserving the span of the matrix’s principal components, allowing leverage score sketches to capture essential information with fewer samples than a Gaussian sketch, especially when $\mathbf{A}$ is incoherent. This point is captured in the case of uniform sampling for an $\ell_{2}$-s.e. in the $5^{th}$ row of Table I, where the embedding dimension is proportional to the coherence of the subspace. In such cases, leverage score sketches can provide accurate results with fewer samples than the required $R$ indicated above, due to the fact that they then perform non-uniform targeted sampling, which often reduces the effective sample size required in practice. The Gaussian and Rademacher sketches which are OSEs drop this dependency on $\gamma(\mathbf{A})$, thus uniform sampling suffices, and their analyses do not need to consider a worst case scenario.
@@ IV Distributed Local Sketching
In this section, we discuss the details of our distributed sketching scheme (Algorithm 1). In the setup of Fig. 1, the $i^{th}$ node applies a random projection matrix $\mathbf{P}_{i}\in\mathbb{R}^{n\times n}$ which is generated locally to flatten the corresponding leverage scores. As we will see, the flattening here is with respect to (w.r.t.) $\mathbf{U}_{i}$; i.e. w.h.p. $\ell_{j}\approx\mathscr{L}_{i}/(nd)$ for each $j\in\mathcal{I}_{i}$. This is the cost we pay for performing local sketching. Nonetheless, we show in Fig. 2 that the flattening degrades gracefully as $k$ increases, even for global datasets with highly non-uniform leverage scores. To partially circumvent this concern if our approach is to be performed by a single user or centrally administered by the coordinator, a random permutation can be applied on the rows of $\mathbf{A}$ before the partitioning takes place.
After locally applying $\mathbf{P}_{i}$, the nodes randomly sample $r=R/k$ rows from $\mathbf{P}_{i}\mathbf{A}_{i}$ which they rescale by $\sqrt{n/r}=\sqrt{N/R}$ and aggregate through $\mathbf{\Omega}_{i}$, to obtain the local sketches
@@
Then, each node communicates the resulting sketch to the coordinator, who aggregates them.
Input: Effective local dimension $r$ $\triangleright$ $R>d$ and $R=rk$
@@
  1)\ 1) Generate a random $\mathbf{P}_{i}\in\mathbb{R}^{n\times n}$ $\triangleright$ $\mathbb{E}\left[\mathbf{P}_{i}^{\top}\mathbf{P}_{i}\right]=\mathbf{I}_{n}$
  2)\ 2) Uniformly sample $r$ rows from $\mathbf{P}_{i}\mathbf{A}_{i}$, through $\mathbf{\Omega}_{i}$
  3)\ 3) Deliver $\sqrt{\frac{n}{r}}\left(\mathbf{\Omega}_{i}\mathbf{P}_{i}\right)\mathbf{A}_{i}\eqqcolon\mathbf{S}_{i}\mathbf{A}_{i}$ to the coordinator
 note: $1)$ and $2)$ can be performed simultaneously by generating $\mathbf{S}_{i}\in\mathbb{R}^{r\times n}$, to reduce the local computations
@@ IV-A Analysis of our approach
[…]
@@
where $\hat{\mathbf{U}}_{i}=\big(\sqrt{n/r}\cdot\mathbf{P}_{i}\big)\mathbf{U}_{i}$ for each $i\in\mathbb{N}_{k}$. We denote the leverage scores of the projected matrix $\hat{\mathbf{P}}\mathbf{A}$ by $\hat{\ell}_{j}=\|\hat{\mathbf{U}}_{(j)}\|_{2}^{2}$.
Next, we show that the leverage scores of each $\hat{\mathbf{U}}_{i}$ are flattened w.r.t. $\mathscr{L}_{i}$ by the local projection. For our analysis, we assume that $\mathbf{P}_{i}$ is a random unitary matrix, drawn from an arbitrary large finite subset $\tilde{O}_{n}(\mathbb{R})$ of the set orthonormal matrices $O_{n}(\mathbb{R})$ of size $n\times n$. Such sketching techniques, in which we first project with a unitary matrix and and then uniformly sample rows from the transformed matrix, are referred to as “Randomized Orthonormal Systems” [49]. The SRHT is a special case and the most popular, due to to the fact the Hadamard matrix is well structured and can be applied efficiently. In practice, normalized Gaussian and Rademacher matrices are also widely used, as they satisfy $\mathbb{E}\left[\mathbf{G}^{\top}\mathbf{G}\right]=\mathbf{I}_{N}$. Analogous results can also be derived for these options of $\mathbf{P}_{i}$.
@@
For a fixed $i\in\mathbb{N}_{k}$ and $\xi>0$, the normalized (w.r.t. $\hat{\mathbf{U}}$) leverage scores $\{\bar{\ell}_{j}\}_{j\in\mathcal{I}_{i}}$ corresponding to $\hat{\mathbf{U}}_{i}$ satisfy
@@
In Fig. 2, we show numerically the flattening of the normalized leverage scores of a random $\mathbf{A}\in\mathbb{R}^{2000\times 40}$ following a $t$-distribution, which scores were highly non-uniform. In this experiments, $\mathbf{P}_{\{k\}}$ were random Gaussian matrices. As noted previously, random Gaussian matrices are good surrogates for unitary random matrices, and are widely used in practice as they are approximately orthogonal. Furthermore, we observe that the flattening degrades gracefully as $k$ increases, and if a random permutation on $\mathbf{A}$’s rows is applied before $\hat{\mathbf{P}}$, we have slightly better results for each $k$. Analogous simulation results were observed when the $\mathbf{P}_{i}$’s were randomized Hadamard transforms, random unitary, and Rademacher random matrices.
The choice for presenting the flattening of leverage scores when $\mathbf{A}\sim t$-distribution is to convey that this occurs even when the scores are highly non-uniform, which is central to our approach. Since flattening occurs w.h.p. when the data points are inhomogeneous, we expect this to also be true with an even higher probability when they are homogeneous. Further experimental verification, considering other synthetic data matrices, can be found in Appendix C.
@@
Before we present our main result, we need to show that $\mathbf{\Omega}$ is close to being a uniform sampling (w.r.) matrix of $R$ out of $N$ rows. In Proposition 2, we show this by applying Chebyshev’s inequality to the balls into bins problem. Specifically, it shows that w.h.p. the sampling of $\mathcal{S}$ in $\mathcal{I}$ is close to a sampling of $\bigcup_{i=1}^{k}\mathcal{S}_{i}$, i.e. the cardinality of sampled indices of $\mathcal{S}$ that lie in any $\mathcal{I}_{i}$ is not far from $\mathcal{S}_{i}=r$. The important factors to note are that all sampling trials in both scenarios are uniform, identical, independent and with replacement.
@@
Partition the sampled index set $\mathcal{S}$ into ordered partitions $S_{i}$ of $\mathcal{I}$ according to $\mathcal{I}_{\{k\}}$, i.e. $S_{i}=\mathcal{S}\bigcap\big(\bigcup_{l=1}^{R}\mathcal{I}_{i}\big)$. Then, for any $i\in\mathbb{N}_{k}$: $\Pr\big[|\#S_{i}-r|\geqslant 10\big]\leqslant 1/100$.
@@
Let $\mathbf{P}_{\{k\}}$ of Algorithm 1 be random unitary matrices, and $\hat{\beta}=\frac{k}{d}\cdot\min_{i\in\mathbb{N}_{k}}\left\{\mathscr{L}_{i}\right\}$. Then, for $\delta>0$ and $R=\Theta\left(d\log{(2d/\delta)}/(\hat{\beta}\epsilon^{2})\right)$, the sketching matrix $\widehat{\mathbf{S}}$ of the global $\mathbf{A}$, satisfies (5) with probability at least $1-\Theta(\delta)$.
@@
[…]